2. What mass of ammonia would be produced by the reaction of 5.0 g of hydrogen?


1
Expert's answer
2021-01-21T03:50:01-0500

N2 +3H2= 2NH3

n(H2)= 5.0/2=2.5 mole

n(NH3)=2.5*2/3= 1.67 mole

m(NH3)= 1.67*(14+3)= 28.39 g

Answer: mass of ammonia is 28.39 g